OpenStudy (anonymous):

Find the inverse of f(x)=6x +7

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

Ok, this is better than looking at a picture. Rewrite as y = 6x + 7. As I said before, when finding inverses, swap the y and x variables. Solve for y again.

OpenStudy (nincompoop):

how about this, solve for x and then switch the x and y at the end

OpenStudy (anonymous):

What will it be

OpenStudy (ipwnbunnies):

You can do it!

OpenStudy (anonymous):

6x-7

OpenStudy (ipwnbunnies):

Go back to y = 6x + 7. Do what nincompoop said, solve for x first.

OpenStudy (anonymous):

Y-7/6

OpenStudy (ipwnbunnies):

Good, so (y-7)/6 = x. Now, just swap x and y, and that'll be the inverse of f(x).

OpenStudy (ipwnbunnies):

Replace y with f(x) also afterwards.

OpenStudy (anonymous):

X-7/6=y

OpenStudy (ipwnbunnies):

\[f ^{-1}(x)\]* I mean

OpenStudy (ipwnbunnies):

Yes, y = (x-7)/6. Replace y with the f^(-1)(x) operator (Idk what it's called).

OpenStudy (nincompoop):

inverse function

OpenStudy (ipwnbunnies):

Idk what you call the actual sign thing, lol. Like 'f(x)'

OpenStudy (ipwnbunnies):

Maybe I'm just thinking too hard, and you just call it function of x. O: